OpenStudy (anonymous):

log(x + 5) - logx = 3

OpenStudy (anonymous):

log((x+5)/x)=3

OpenStudy (anonymous):

One of my multiple choice answers is 2.5, would that work? It's the closet one to 3..lol

OpenStudy (anonymous):

No 3 is not the answer @vt you have to solve the equation

OpenStudy (anonymous):

Oh. Lol that's unfortunate. I don't have a scientific calculator.

OpenStudy (anonymous):

use google

OpenStudy (anonymous):

They have one?

OpenStudy (anonymous):

yes! you will find an online calculator

OpenStudy (anonymous):

Alright. Let's see if I can figure it out lol.

OpenStudy (isaiah.feynman):

Remember log rules??

OpenStudy (anonymous):

No.. I really dont have a clue tbh haha

OpenStudy (isaiah.feynman):

This might be a little tedious for us..

OpenStudy (anonymous):

Lol just a little.

OpenStudy (isaiah.feynman):

Since is a difference of two logs, we use the rule log a - log b = log(a/b)..got it?

OpenStudy (anonymous):

Alright, so far so good.

OpenStudy (isaiah.feynman):

So your problem becomes..

OpenStudy (isaiah.feynman):

That's what your problem becomes

OpenStudy (anonymous):

I would probably be able to solve it on a calculator but I just don't know where to find x on the calculator....lol

OpenStudy (isaiah.feynman):

No. I don't recommend that. This is straight forward.

OpenStudy (anonymous):

alright alright i feel u. keep on keepin on then

OpenStudy (isaiah.feynman):

Now we see that the logarithm has no obvious base... but by convention we take the base to be 10. Then using another rule we transform from logarithmic to exponential form.

OpenStudy (isaiah.feynman):

Get it?

OpenStudy (anonymous):

Yessir

OpenStudy (isaiah.feynman):

So can you solve for x?

OpenStudy (radar):

What are your options as far as choices go?

OpenStudy (radar):

I get x as approximately 5 10^-3 or .005

OpenStudy (radar):

I was using what @vt linked to. X= 5/999 approx 5/1000 or .005 or 5 10^-3

OpenStudy (radar):

Excuse me @Isaiah.Feynman linked to.(not vt)

OpenStudy (anonymous):

5/999?

OpenStudy (anonymous):

.005 yep yep :) thanks everyone!